OpenStudy (anonymous):

The sum of three consecutive odd integers is –27. What are the numbers?

OpenStudy (solomonzelman):

-27 / 3 = -9 -10 -9 -8

OpenStudy (solomonzelman):

My bad -11 -9 -7

OpenStudy (solomonzelman):

it asks odd, I found the average number.

OpenStudy (zpupster):

here is another way to solve: n-1= smallest odd integer n-3= next odd and n-5 is next odd n-1+N-3+n-5= -27 3n=-18 n=-6 then -6-1=-7 etc